OpenStudy (anonymous):

Whis 6/12 divided by 6/13

Nnesha (nnesha):

change division to multiplication multiply the `numerator` with the ` reciprocal` of the bottom fraction

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ \frac{ a }{ b }}{ \frac{ c }{ d }}=\frac{ a }{ b} \times \frac{ d }{ c}\]